Problems with PS Tricks

Christian Kral August 13, 2010

Abstract

1 Here comes the problem

The DVI file cannot be converted using ps2pdf13 when input is used.

Zweig 1 Zweig 2 1 2 3 4 5 6 7 8 9 10 11 121 36 6 12 18 24 30

Figure 1: This is the PS Tricks example